# Intensity-Efficient and Intensity-Positional Allocations

### 01 Mar 2026 v6

@@ Abstract
This paper studies the allocation problem in environments where, in addition to having ordinal preferences, agents also have ordinal intensities: they can say “I prefer $a$ to $b$ more than I prefer $c$ to $d$ without necessarily being able to claim how much more. A rank-based criterion for interpersonal comparisons of ordinal intensities is introduced for this analytical environment. Building on it, the concepts of intensity-dominated, intensity-efficient and intensity-positional allocations are defined.Two allocations are related by intensity dominance if they assign the same two items to the same two agents but in a “flipped” way, and one allocation assigns the commonly preferred item in any such pair to the agent who prefers it more. An allocation is intensity-efficient if it is Pareto-efficient and not intensity-dominated. Distinctly, an allocation is intensity-positional if it maximizes the agents’ total intensity-rank scores. The latter extend standard Borda scores to environments where consecutively-ranked items do not necessarily feature identical intensity differences. The paper analyses intensity-efficient and intensity-positional allocations, and clarifies how they differ from those based on the utilitarian or Borda rules.
@@ 1 Introduction
Distributional principles that are used in the assignment of goods to individuals can be evaluated jointly on the realism of their assumptions and the appeal—according to some justice criterion—of their prescribed allocations. Arguably the most important such economic principle is Pareto-efficiency: assuming only information on agents’ preference rankings over the relevant goods, its notion of optimality is defined by the requirement that no changes can make some individuals better off without hurting others. Yet it is well-known that among the typically many Pareto-efficient allocations there are often some which violate intuitive notions of distributive justice. Appreciation of these facts raises the question of whether one can narrow down and improve upon Pareto’s criterion by considering alternative principles which, in addition to agents’ preferences, also incorporate information on their preference intensities.
Introducing and analyzing such a principle, in turn, requires clarity on what is meant by preference intensities. Most of economic theory has traditionally taken the existence of intensity comparisons as synonymous to the existence of a cardinal utility function, and has treated the way in which utility differences derived from such a function are ranked as the determinant of how the individual’s intensity comparisons are ordered. Building on cardinal utility functions over goods that are derived from expected-utility preferences over lotteries over goods (von Neumann and Morgenstern, 1947), and assuming that utility differences are interpersonally comparable, the classical utilitarian summation criterion, formalized in Harsanyi (1955), is the benchmark distributional principle that refines the set of Pareto-efficient allocations in this analytical environment (Moulin, 1988, 2004).
Despite the undeniable tractability and widespread application of this and other utilitarian criteria in various domains of economic analysis, it is well-understood that the specific cardinal utility functions emerging from expected-utility preferences over lotteries over goods confound an individual’s intensity comparisons with that person’s attitudes toward risk (Arrow, 1951, 1963; Luce and Raiffa, 1957; Baccelli and Mongin, 2016; Sen, 2017; Alaoui and Penta, 2025; Samuelson, 2025). More importantly, when cardinal utility differences —even with riskless axiomatic primitives—are used to define an agent’s intensity comparisons, the agent in question is portrayed as having a superhumanly high degree of precision in their capacity to quantify those intensities (Samuelson, 1938; Basu, 1982). This stands in contrast to some contemporary views in the field of cognitive psychology and decision-making, according to which “the perceptual system might be like a pan in balance, which responds by tipping to the left or right, depending on which of two items is heavier, but provides no read-out of the absolute weight of either item.” (Vlaev et al., 2011, p. 551). To the extent that comparing options is easier and more reliable than assigning values to them, it is reasonable to hypothesize that comparing preference-intensity differences in a similar fashion is also cognitively less demanding than assigning numerical values to them, hence possibly a more accurate reflection of the individual’s true intensity comparisons.
Motivated by this tension between the intuitive appeal of including intensity information in the allocation process and what might be regarded as weaknesses of the standard cardinal-welfarist approach toward doing so, this study introduces and analyses a novel third domain for social-choice and matching-theoretic analysis which provides a middle ground for further progress to be made. Specifically, instead of cardinal-welfarist or Arrowian profiles that, respectively consist of interpersonally comparable cardinal utility functions—possibly not arising from expected-utility preferences (Nebel, 2024; Weymark, 2005; Harvey, 1999)—or preference orderings, this domain comprises profiles of generally ordinal preference intensity orderings, i.e. collections of quaternary relations that reflect the agents’ preference intensity comparisons. This analytical choice assumes that, similar to the way in which preferences are elicited, e.g. in matching problems, information about agents’ intensities is obtainable by asking them to respond to simple comparison questions such as “Do you prefer $Apples$ to $Berries$ more than you prefer $Pears$ to GrapesGrapes?”, and that these responses are internally consistent in a way that we make precise in Section 2.
This framework where both preference and—generally only ordinal—intensity information is available to the social planner naturally raises the question of how it might be used, as per the goals that we set out above, to arrive at some normatively appealing refinement of Pareto-efficiency that would also reflect differences in the agents’ intensities and mitigate the multiplicity of Pareto-efficient allocations. Similar to utilitarian and related cardinal-welfarist criteria, such a refinement requires some kind of interpersonal comparisons. Unlike those notions, however, where such comparisons are based on agents’ utilities (Fleurbaey and Hammond, 2004; d’Aspremont, 2018; Echenique et al., 2023; Baccelli, 2023), in our framework they must rely instead on the information contained in the above ordinal intensity rankings.
To make progress against this analytical challenge, and focusing on the assignment problem with finitely many indivisible goods, we submit that such comparisons can be made by contrasting the rank-order position of pairs of alternatives $(a,b)$ in the different agents’ intensity orderings, whenever such contrasting is meaningful. In particular, when both agents $i$ and $j$ prefer $a$ to $b$ but the pair $(a,b)$ lies higher in $i$’s intensity ranking than in $j$’s, and these rankings have the same length, then we interpret this as suggesting that $i$ prefers it more. In Section 3 we formalize and motivate this interpretational assumption concerning interpersonal comparisons of ordinal intensities in the present analytical environment. We also clarify that, unlike utilitarian environments where interpersonal comparisons are driven by agents’ absolute intensities that are captured by the size of the respective utility differences, this postulate instead is rooted in comparing agents’ relative intensities in a scoring-like fashion that is analogous in this environment to the well-known Borda-count (Borda, 1781; Young, 1974; Mas-Colell et al., 1995; Maskin, 2025) or relative-utilitarian (Dhillon and Mertens, 1999) normalizations in ordinal- and cardinal-preference environments, respectively (d’Aspremont, 2018; Baccelli, 2023).
In Section 4 we build on this comparability postulate and introduce a refinement of Pareto efficiency that incorporates such interpersonal comparisons of ordinal intensities, with the goal of arriving at allocations that are more distributively just. Specifically, we define an allocation $x$ to be intensity-efficient if it is Pareto-efficient and also such that, whenever another Pareto-efficient allocation $y$ assigns the same pairs of objects to the same pairs of agents but in a “flipped” way, i.e. when $(x_{i},x_{j})=(y_{j},y_{i})=(a,b)$ for agents $i$, $j$ and alternatives $a$, $b$, then $x$ assigns the commonly preferred alternative in every such pair to the agent who prefers it more. Assuming strict preferences and intensities, we first show that an intensity-efficient allocation exists under every profile when there are three agents and alternatives. We also show that existence is not guaranteed when there are more agents and items, because the underlying dominance relation may be then be cyclic.
@@
Motivated by this, for the general case of $n$ agents and as many items we establish a sufficient condition for existence, requiring profiles to be (effectively) homeo-monotonic. The key part of this condition could be summarized informally as follows: when a group of $k$ agents have identical preferences over their top $k$ items, and agents can be weakly ordered in how much they prefer any item in this subset over that item’s immediate successor, then this ordering cannot be violated when it comes to comparing the agents’ intensities between any non-consecutively ranked items in this subset. Importantly, although the presence of such groups—and the corresponding restrictions within them—is not necessary for existence, we show that if a profile does give rise to such a group, and if the above-mentioned ordering over its agents is strict, then homeo-monotonicity in that group is indeed necessary for an intensity-efficient allocation to exist.
@@
Intensity-efficiency appears to be the first intensity-inclusive refinement of Pareto-efficiency that is operational in an environment where neither the agents’ utilities are required to be intra- or inter-personally comparable nor monetary transfers between agents are assumed to be feasible. Yet an interesting by-product of the existence investigation is the realization that, in those analytical environments where all agents’ intensity comparisons do emerge from cardinal utility functions that are assumed to be interpersonally comparable, classical-utilitarian and intensity-efficient allocations do not generally coincide. They do so, however, when the said utility profile satisfies an intuitive additional condition that we introduce. This requires any interpersonal dominance that emerges by comparing two agents’ utility differences to be accompanied by an interpersonal dominance in the same direction in the rank-order positions of these utility differences in the respective agents’ intensity orderings.
@@
In Section 5 we build on a stronger interpersonal comparisons assumption to propose and analyse the concept of intensity-positional allocations. Toward defining them, we recall first than an item’s Borda score for some agent is the number of items that are ranked below it by that agent. Although Borda scores give rise to a specific normalized ordinal utility representation of the agent’s preferences, in practice they are often used with a cardinal interpretation (see, for example, Maskin, 2025) due to the fact that ordering the utility differences induced by this normalization defines the special class of intensity relations where the agent prefers two alternatives that are ranked $k$ places apart exactly as much as she prefers any two other alternatives that are ranked $k$ places apart. We will refer to these intensity relations as “linear”. Similar to Borda allocations, which are defined as those that maximize the aggregate Borda score, intensity-positional allocations are defined as the maximisers of the aggregate intensity-rank score. Unlike an item’s Borda score, however, which reveals the position of that item on the agent’s preference ordering, its intensity-rank score corresponds to the highest rank-order position in the agent’s intensity ordering where this item appears as preferred to something else. Hence, it encodes information both about the number of alternatives that are ranked below this item and about the relative intensity differences between it and the items it is preferred to. This information, in turn, is reflected in the prescribed allocation process. In the paper’s focal case of strict intensity profiles we show that intensity-positional allocations refine standard Borda allocations. We further show that, in a larger domain where a like-for-like comparison remains meaningful, intensity-rank and Borda-score aggregation diverge in general but coincide at linear profiles. The intuitive reason is that, unlike intensity-positional aggregation, Borda aggregation treats every profile as it if were linear. Thus, the intensity-positional criterion is a genuine generalization of the standard Borda-count criterion in this extended domain.
Intensity-positional and intensity-efficient allocations are logically distinct. Two advantages of the former relative to the latter are their generic existence and their incorporation of a way to resolve intensity-pertinent welfare trade-offs across agents. A disadvantage is that making intensity-positional allocations operational implicitly invokes stronger interpersonal comparability assumptions.
In Section 6, finally, we turn to the problem of designing a matching mechanism that assigns an intensity profile to an allocation that improves upon the Pareto-efficiency desideratum. We first show that once agents’ ordinal preferences have been elicited, their ordinal intensities that induce those preferences can be elicited in quadratic time by leveraging the monotonic “lateral consistency” relation that exists between preferences and intensities in the class that we focus on. Then, building on the matching-theoretic literature with one-sided preferences, we study an algorithm that extends the fundamental Random Priority (RP) mechanism’s algorithm and produces—in quadratic time in the number of items/agents—an allocation which intensity-dominates the Pareto-efficient one that is found by RP. While RP makes the elicitation of agents’ preferences incentive-compatible, we show that this does not carry over to the elicitation of agents’ intensities under the proposed extension of the algorithm. We connect our analysis to related recent work in this literature that started with Abdulkadiroğlu, Che, and Yasuda (2011) and focuses on the trade-off between utilitarian welfare and incentive-compatibility that is relevant in a designer’s choice between ordinal and cardinal matching mechanisms.
@@
As far as the broader related literature is concerned, this paper’s contribution lies at the intersection of welfare economics, social choice and matching theory, in all three cases building bridges between the existing ordinal- and cardinal-welfarist paradigms. Similar to existing cardinal-welfarist refinements of Pareto-efficiency such as those based on classical or relative utilitarianism (Harsanyi, 1955; Dhillon and Mertens, 1999), Nash social welfare (Nash, 1950; Luce and Raiffa, 1957; Moulin, 2004) and others, the proposed refinement is inclusive of information that goes beyond agents’ preferences toward also reflecting their intensities in the allocation process. Unlike these refinements, the hereby proposed ones are the first that are operational in the general class of ordinal preference-intensity profiles, and build on novel interpersonal-comparability criteria that are distinct from interpersonal comparisons of utility differences which the above notions rely on. Also unlike these refinements, however, intensity-efficient allocations are introduced and studied here directly in the context of the assignment problem, rather than as the application of some social welfare function that is generally defined on the relevant domain. Intensity-positional allocations on the other hand are indeed derived from a social welfare function. The problem of characterizing this function axiomatically is important and deserves separate treatment.
@@
In addition to Abdulkadiroğlu, Che, and Yasuda (2011), finally, several studies in matching theory—both with one- and two-sided preferences—have taken agents’ interpersonally comparable von Neumann-Morgenstern cardinal utilities (Budish, 2011; Lee and Yariv, 2018; Ortoleva, Safonov, and Yariv, 2023) or quasi-linear preferences in willingness to pay (Che, Gale, and Kim, 2013) as inputs into the allocation process toward maximizing utilitarian welfare. Our study shares the same goal as those above, namely to improve upon the distributive-justice properties of Pareto-efficiency in the final matching, yet pursues it in the unexplored environment where only ordinal intensity comparisons are generally available to the designer. While this environment is less tractable for the modelling analyst, it is simpler for the economic agent and, as such, we believe worthy of further investigation.
2 Measurement-Theoretic Foundations

### 14 Apr 2026 v8

# Efficient Allocation with Ordinal Preference Intensities

@@ Abstract
Standard ordinal allocation methods ignore how strongly agents value different improvements, while cardinal methods require additional assumptions that are often considered too demanding. This paper studies assignment problems in the middle ground environment of ordinal preference intensities where agents can rank alternatives as well as preference improvements. The two criteria it proposes—intensity-efficient and intensity-positional allocations—use this combined information to refine Pareto efficiency via a novel dominance concept and a generalization of Borda-style scoring, respectively. These criteria point in new directions where stronger welfare gains are possible without cardinal utility or monetary transfers assumptions.
@@ 1.1 Motivation
Distributional criteria for the assignment of goods to individuals can be evaluated jointly on the realism of their assumptions and the appeal—according to some justice principle—of their prescribed allocations. Arguably the most important such criterion is Pareto efficiency: assuming only information on agents’ preferences over the relevant goods, its notion of optimality is defined by the requirement that no changes can make some individuals better off without hurting others. Yet it is well-known that among the typically many Pareto-efficient allocations there are often some which violate intuitive notions of distributive justice. Appreciation of these facts raises the question of whether Pareto’s criterion might be improved on by considering alternative principles which, in addition to agents’ preferences, also incorporate information on their preference intensities. Introducing and analyzing such a principle, in turn, requires clarity on what is meant by preference intensities. Economists have traditionally interpreted these to be synonymous to the existence of a cardinal utility function that represents an agent’s preferences over alternatives. Typically, this has been of the von Neumann and Morgenstern (1947) (vNM) kind, derived from expected-utility preferences over lotteries over alternatives. Assuming that utility differences from such functions are interpersonally comparable, the classical utilitarian summation criterion, formalized in Harsanyi (1955), has been the benchmark distributional principle of the cardinal-welfarist tradition (Moulin, 1988; Roemer, 1996; Fleurbaey and Hammond, 2004; d’Aspremont, 2018; Echenique, Immorlica, and Vazirani, 2023), refining the set of Pareto-efficient allocations in intensity-inclusive directions.
Despite the undeniable tractability and widespread application of this and related criteria in various domains of economic analysis, it is well-understood that cardinal utility functions emerging from expected-utility preferences over lotteries confound an individual’s intensity comparisons with that person’s attitudes toward risk (Arrow, 1951, 1963; Luce and Raiffa, 1957; Baccelli and Mongin, 2016; Sen, 2017; Alaoui and Penta, 2025; Samuelson, 2025). This fact, in turn, has generated concerns—such as the one expressed by Kenneth Arrow in the cited quotation—about the epistemic origins of such functions. In addition, and no less importantly, when cardinal utility differences are used to define an agent’s intensity comparisons, the person in question is portrayed as having a supernaturally high degree of precision in their capacity to quantify those intensities (Samuelson, 1938; Basu, 1982). Acknowledging the logical implication of this fact—namely, that elicitation of agents’ utilities in reality may be subject to “noise”—diminishes a planner’s confidence that application of utilitarian or related criteria in practical allocation problems is guaranteed to achieve the intended distributive outcome. In the distinct but related domain of evaluative voting problems, for example, where agents rate the candidates on numerical scales, Baujard, Gavrel, Igersheim, Laslier, and Lebona (2018); Baujard, Igersheim, and Lebon (2021) provide experimental evidence suggesting that the use of different scales leads to non-equivalent voting results.
1.2 Contribution
This skepticism notwithstanding, intensity comparisons such as those in Harsanyi’s (1987) opening quotation are common in everyday life and, in fact, do not presuppose the existence of interpersonally comparable utilities. They do, however, implicitly assume that:
(i) agents can rank alternatives according to their preferences and rank preference improvements according to the intensity of these preferences;
(ii) systematic interpersonal comparisons of such second-order data are possible.
Given all the above, this paper’s central question can now be stated: “When agents cannot provide credible cardinal utilities, but can report which preference improvements feel larger than others, can a planner use that information to improve upon Pareto-efficient allocations?” We answer this question by focusing on the classic assignment problem where finitely many indivisible goods must be assigned to as many agents.
To address requirement (i) we assume that agents’ preference intensity orderings belong to the broad class of those that admit a generally ordinal numerical representation by means of a preference intensity function (Gerasimou, 2021). In effect, this portrays each individual as being able to rank the above-mentioned preference improvements in a coherent way, yet without necessarily being able to quantify them. This modeling approach is one possible way of dealing with the concern expressed earlier regarding cardinal utilities, and is in line with some contemporary views in cognitive psychology that are elegantly summarized in Vlaev, Chater, Stewart, and Brown (2011, p. 551) with the remark that “the perceptual system might be like a pan in balance, which responds by tipping to the left or right, depending on which of two items is heavier, but provides no read-out of the absolute weight of either item.” To address requirement (ii), moreover, we employ an intuitive normalization of agents’ preference intensity functions and introduce an interpersonal comparability assumption that is suitable for this kind of ordinal intensity data. With this interpretational assumption in the background, finally, we propose two refinements of Pareto efficiency—intensity-efficiency and intensity-positionality—that incorporate such data and discard allocations that are intuitively inferior to others in the light of this additional information.
To illustrate both the interpersonal comparability assumption and the intensity-efficiency criterion, imagine, for example, that Anna and Baba have identical preferences that put ($a$)pple above ($b$)anana, and the latter above ($o$)range. Suppose Anna thinks the preference improvement from b to a is bigger than the improvement from o to b, and that the opposite is true for Baba (clearly, the improvement from o to a is the biggest one for both). Who should get the apple if it is also known that the third person in the group likes the orange most? The Pareto criterion, as well as the Pareto-refining Borda criterion (Borda, 1781; Young, 1974; Mas-Colell et al., 1995) that we recall two paragraphs below, delivers a tied recommendation based solely on the agents’ preferences: either Anna or Baba should get it. By contrast, the intensity-efficiency criterion assigns it to Anna. The reason is that the preference improvement from b to a lies in a higher position in Anna’s intensity ordering than in Baba’s. The interpersonal comparability condition translates this fact into Borda-style intensity scores on a common scale (ranging from 1 to 3 here) and, in that spirit, declares Anna’s improvement to be stronger than Baba’s because pair $(a,b)$ has a score of 2 for Anna and 1 for Baba. Intensity-efficiency requires that, other things equal, whenever two Pareto allocations assign the same two items to the same two agents but in a “flipped” way, and one allocation assigns the commonly preferred item in every such pair to the agent who prefers it more, then that allocation should be favored over the alternative, which is discarded as being intensity-dominated.
@@ 1.2 Contribution
Assuming strict preferences and ordinal intensities, in Theorem 1 we show that an intensity-efficient allocation exists under every profile when there are three agents and alternatives. Existence is not guaranteed when there are more because the underlying intensity-dominance relation may be cyclic. For the general case of $n$ agents and items, then, we establish a sufficient condition for existence, requiring profiles to be (effectively) homeo-monotonic (Theorem 2). The key part of this condition can be summarized informally as follows: when a group of $k\leq n$ agents have identical preferences over their top $k$ items, and the $k$ agents can be weakly ordered in how much they prefer any item in this subset over that item’s immediate successor, this ordering cannot be violated when agents’ intensities between any non-consecutively ranked items are compared. For example, consider a group that consists of agents $L_{1}$, $L_{2}$, $L_{3}$ and items $l_{1}$, $l_{2}$, $l_{3}$ over which agents have the identical preferences $l_{1}\succ l_{2}\succ l_{3}$ but $L_{1}$ ranks the improvements from $l_{3}$ to $l_{2}$ and from $l_{2}$ to $l_{1}$ higher than $L_{2}$, who ranks them higher than $L_{3}$. Then, homeo-monotonicity requires the improvement from $l_{3}$ to $l_{1}$ be ranked weakly higher by $L_{1}$ too, followed by $L_{2}$ and then by $L_{3}$. Although the presence of such groups is not necessary for intensity-efficient allocations to exist, if a profile gives rise to a group with the antecedent structure over consecutively-ranked items, and if the relevant ordering over its agents is strict, then homeo-monotonicity within that group is indeed necessary for existence (Theorem 3).
@@
Intensity-efficiency is the first intensity-inclusive refinement of Pareto-efficiency that operates in an environment where neither the agents’ utilities are required to be intra- or inter-personally comparable nor monetary transfers between agents are assumed to be feasible. Yet an interesting by-product of the existence investigation is the realization that, in those analytical environments where all agents’ intensity comparisons do emerge from cardinal utility functions that are interpersonally comparable, classical-utilitarian and intensity-efficient allocations do not generally coincide (Proposition 1). The reason is that an agent’s utility difference between two items may be greater than another agent’s in absolute terms but still have a lower intensity score. When no such discrepancy arises, utilitarian allocations are indeed intensity-efficient (Proposition 2).
@@
By its definition, the intensity-dominance relation on which the intensity-efficiency criterion rests is not always applicable. Motivated by this fact, we later introduce a stronger interpersonal comparability assumption and, building on it, propose the concept of intensity-positional allocations. Toward defining it, recall that an item’s Borda score for some agent is the number of items that this person ranks below it. Importantly, although Borda scores give rise to a specific normalized ordinal utility representation of the agent’s preferences, in practice these are often used with a cardinal interpretation (see, for example, Maskin, 2025 and, for recent experimental evidence, Ambuehl and Bernheim, 2026). This is due to the fact that ordering the utility differences induced by this normalization defines the special class of intensity relations where the agent prefers two alternatives that are ranked $k$ places apart exactly as much as she prefers any two other alternatives that are ranked $k$ places apart. We will refer to these intensity relations as “linear”. Similar to Borda allocations, which are defined by maximization of the aggregate Borda score, intensity-positional allocations are those that maximize the aggregate intensity-rank score. Unlike, say, item $a$’s Borda score, however, which reveals the relative position of $a$ in the agent’s preference ordering, $a$’s intensity-rank score corresponds to the highest-ranked preference improvement towards item $a$. The latter, in turn, is simply the position of pair $(a,w)$ in the agent’s intensity ordering, where $w$ is that person’s bottom-ranked item.
An item’s intensity-rank score, therefore, encodes information both about the number of alternatives that are ranked below it and about the relative intensity differences between it and the items it is preferred to. This information, in turn, is reflected in the prescribed allocation process. In the paper’s focal case of strict intensity profiles we show that intensity-positional allocations refine Pareto-efficient allocations and are distinct from Borda allocations (Propositions 3 and 4). We further show that, in a larger domain where a like-for-like comparison remains meaningful, intensity-rank and Borda-score aggregation diverge in general but coincide at linear profiles (Theorem 4). The intuitive reason is that, unlike intensity-positional aggregation, Borda aggregation treats every profile as if it were linear. Thus, the intensity-positional criterion is a genuine generalization of the standard Borda-count criterion in this extended domain, enabling the analyst to select Pareto-efficient allocations with desirable welfare properties which might otherwise have been discarded. Intensity-positional and intensity-efficient allocations are logically distinct. Two advantages of the former relative to the latter are their generic existence and their incorporation of a way to resolve intensity-pertinent welfare trade-offs across agents. A disadvantage is that making intensity-positional allocations operational implicitly invokes stronger interpersonal comparability assumptions.
1.3 Related Literature
@@
The paper contributes to welfare economics, social choice and allocation theory, in all three cases building bridges between the ordinal- and cardinal-welfarist paradigms. In particular, similar to cardinal-welfarist refinements of Pareto-efficiency such as those based on classical or relative utilitarianism (Harsanyi, 1955; Dhillon and Mertens, 1999), the hereby proposed refinements of Pareto efficiency are inclusive of information that goes beyond agents’ preferences toward also reflecting their intensities in the allocation process. Unlike the former refinements, those studied here are the first that are applicable in principle in the general class of ordinal preference-intensity profiles. Furthermore, they build on a novel analytical environment and interpersonal-comparability criteria that are distinct from interpersonal comparisons of cardinal utility differences. The previous subsection also clarifies how they differ from the ordinal but “interpretationally cardinalized” classic Borda scoring method.
@@
Although the building blocks differ in important ways, conceptually this paper is closely related to Abdulkadiroğlu, Che, and Yasuda (2011) (see also Miralles, 2009; Pycia, 2011; Carroll, 2018) and the several studies in matching theory with one- and two-sided preferences that have taken agents’ interpersonally comparable vNM cardinal utilities (Budish, 2011; Lee and Yariv, 2018; Ortoleva, Safonov, and Yariv, 2023) or quasi-linear preferences in willingness to pay (Che, Gale, and Kim, 2013) as inputs into the allocation process toward maximizing utilitarian welfare. Our study shares the same goal as those above, namely to improve upon the distributive-justice properties of Pareto-efficiency in the final matching by incorporating intensity-relevant information. Its distinguishing point is that it is the first to pursue this goal in the unexplored environment where only ordinal intensity comparisons are generally available to the designer. While this environment is less tractable for the modelling analyst, it is cognitively simpler for the economic agent and, as such, we believe worthy of further investigation.
2 Preliminaries